# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ H.6 Proof of Lemma 9
On one hand, we have $\nabla_{2}\tilde{G}_{\mu}^{i}\big{(}\mathrm{T}_{s},\mathrm{T}_{s}(x)\big{)}=\nabla\frac{\partial}{\partial x_{i}}\frac{\delta\mathcal{F}}{\delta\mu}\big{(}(\mathrm{T}_{s})_{\#}\mu\big{)}\big{(}\mathrm{T}_{s}(x)\big{)}$. On the other hand, let us compute $\nabla_{1}\tilde{G}_{\mu}^{i}(\mathrm{T},x)$. First, we define the shorthands $\tilde{g}_{\mu}^{x,i}(\mathrm{T})=\tilde{G}_{\mu}^{i}(\mathrm{T},x)=\frac{\partial}{\partial x_{i}}\frac{\delta\mathcal{F}}{\delta\mu}(\mathrm{T}_{\#}\mu)(x)$ and $g^{x,i}(\nu)=\frac{\partial}{\partial x_{i}}\frac{\delta\mathcal{F}}{\delta\mu}(\nu)(x)$. Since $\tilde{g}_{\mu}^{x,i}(\mathrm{T})=g^{x,i}(\mathrm{T}_{\#}\mu)$, applying Proposition 8, we know that $\nabla_{1}\tilde{G}_{\mu}(\mathrm{T},x)=\nabla\tilde{g}_{\mu}^{x,i}(\mathrm{T})=\nabla_{\mathrm{W}_{2}}g^{x,i}(\mathrm{T}_{\#}\mu)\circ T$.
@@ H.8 Proof of Proposition 11
    $\begin{cases}\langle\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{1}},\mathrm{T}_{t_{2}}-\mathrm{T}_{t_{1}}\rangle_{L^{2}(\mu)}\leq\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}-\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}\\ \langle\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{2}},\mathrm{T}_{t_{1}}-\mathrm{T}_{t_{2}}\rangle_{L^{2}(\mu)}\leq\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}-\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}\\ \end{cases}$ (162)
@@
    $\langle\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{2}}-\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{1}},\mathrm{T}_{t_{2}}-\mathrm{T}_{t_{1}}\rangle_{L^{2}(\mu)}\geq 0$ (163)
@@
    (c4) $\implies$ (c1). Let $t_{1},t_{2}\in[0,1]$ and $\varphi(t)=\mathcal{F}(\tilde{\mu}_{t}^{t_{1}\to t_{2}})$ for all $t\in[0,1]$. From [114, Equation 16.5],
@@
    where $G$ is the Green function defined as $G(s,t)=s(1-t)\mathbbm{1}_{\{s\leq t\}}+t(1-s)\mathbbm{1}_{\{t\leq s\}}\geq 0$ [114, Equation 16.6]. Then, $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t})\geq 0$ implies that $\int_{0}^{1}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\varphi(s)G(s,t)\ \mathrm{d}s\geq 0$, and thus
    $\varphi(t)=\mathcal{F}(\tilde{\mu}_{t}^{t_{1}\to t_{2}})\leq(1-t)\varphi(0)+t\varphi(1)=(1-t)\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}+t\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}.$ (168)
H.9 Proof of Proposition 22
@@
$\nabla\mathrm{J}(\tilde{\mathrm{T}}_{k+1})$ $=\nabla\phi_{\mu_{k}}(\tilde{\mathrm{T}}_{k+1})-\nabla\phi_{\mu_{k}}(\mathrm{Id})+\tau\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{G}(\mu_{k})+\nabla_{\mathrm{W}_{2}}\mathcal{H}\big{(}(\tilde{\mathrm{T}}_{k+1})_{\#}\mu_{k}\big{)}\circ\tilde{\mathrm{T}}_{k+1}\big{)}$ (169)
@@
$=\nabla\phi_{\mu_{k}}(\tilde{\mathrm{T}}_{k+1})+\tau\nabla_{\mathrm{W}_{2}}\mathcal{H}\big{(}(\tilde{\mathrm{T}}_{k+1})_{\#}\mu_{k}\big{)}\circ\tilde{\mathrm{T}}_{k+1}-\nabla\phi_{\mu_{k}}(\mathrm{S}_{k+1}).$
@@
But $\mathrm{T}_{\phi_{\nu_{k+1}}}^{\nu_{k+1},\mu_{k+1}}\in L^{2}(\nu_{k+1})$, so this is a contradiction. So, we can conclude that the two schemes are equivalent, and moreover, $\tilde{\mathrm{T}}_{k+1}=\mathrm{T}_{\phi_{\nu_{k+1}}}^{\nu_{k+1},\mu_{k+1}}\circ\mathrm{S}_{k+1}$.
H.10 Proof of Proposition 23
Let $\psi(\mathrm{T})=\tau\big{(}\langle\nabla_{\mathrm{W}_{2}}\mathcal{G}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\mathcal{H}(\mathrm{T}_{\#}\mu_{k})\big{)}$. Since ~$\tilde{\mathcal{H}}_{\mu_{k}}$ is convex, $\psi$ is convex, and we can apply the three-point inequality (Lemma 27) and for all $\mathrm{T}\in L^{2}(\mu_{k})$,
@@ H.10 Proof of Proposition 23
$\mathcal{H}(\mu_{k+1})+\mathcal{G}(\mu_{k+1})$ $\leq\mathcal{H}(\mu_{k+1})+\mathcal{G}(\mu_{k})+\langle\nabla_{\mathrm{W}_{2}}\mathcal{G}(\mu_{k}),\tilde{\mathrm{T}}_{k+1}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\tilde{\mathrm{T}}_{k+1},\mathrm{Id})$ (179)
@@
$\quad-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\tilde{\mathrm{T}}_{k+1}).$
@@
Plugging this into (179), we get
@@
Using $\mathrm{T}=\mathrm{Id}$ in (179), we observe that $\mathcal{F}(\mu_{k})\leq\mathcal{F}(\mu_{\ell})$ for all $\ell\leq k$. Moreover, $\mathrm{W}_{\phi}(\nu,\mu_{k})\geq 0$. Thus, applying Lemma 28 with $f=\mathcal{F}$, $c=\mathcal{F}(\nu)$ and $g=\mathrm{W}_{\phi}(\nu,\cdot)$, we obtain
H.11 Proof of Lemma 24
@@
and thus with $u$ being $(1-\tau\beta)$-convex relative to $V$ (for $\tau\beta\leq 1)$. For $\tau\beta<1$, it is equivalent with $u-(1-\tau\beta)V$ convex, i.e. $\langle\nabla u(x)-\nabla u(y),x-y\rangle\geq(1-\tau\beta)\langle\nabla V(x)-\nabla V(y),x-y\rangle\geq 0$. Since $V$ is strictly convex, $\nabla u$ is injective.
@@
To conclude, applying [4, Lemma 5.5.3], $\mathrm{T}_{\#}\mu$ is absolutely continuous with respect to the Lebesgue measure.
H.12 Proof of Proposition 25
@@
Recall from (62) that $\mathrm{T}_{k+1}(x)=\big{(}(1-\tau)\Sigma_{k+1}\Sigma_{k}^{-1}+\tau\Sigma_{k+1}\Sigma^{-1}\big{)}x+cst$, thus $\nabla\mathrm{T}_{k+1}$ is a constant. Using the computations of [40, Appendix B.2],
@@
Moreover, by [31, Lemma 10], $\mu\mapsto\|\Sigma_{\mu}\|_{\mathrm{op}}$ is convex along generalized geodesics, and thus $\Sigma_{\mu_{t}}\preceq\lambda I_{d}$ and $\|\Sigma_{\mu_{t}}\|_{\mathrm{op}}\leq\lambda$ [40]. Hence, noting $\sigma_{\mathrm{max}}(M)$ the largest eigenvalue of some matrix $M$,
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{H}(\mu_{t})\geq\frac{1}{\lambda}\|\mathrm{T}_{k+1}-\mathrm{Id}\|_{L^{2}(\mu_{k})}^{2}$ $\geq\frac{1}{\lambda\sigma_{\mathrm{max}}(\nabla^{2}V)}\int(\mathrm{T}_{k+1}-\mathrm{Id})^{T}\nabla^{2}V(T_{k+1}-\mathrm{Id})\mathrm{d}\mu_{k}$ (192)
$=\frac{2}{\lambda\sigma_{\mathrm{max}}(\nabla^{2}V)}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{V}(\mu_{t}).$
@@
$\mathrm{d}_{\tilde{\mathcal{F}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})$ $=\mathrm{d}_{\tilde{\mathcal{V}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})+\mathrm{d}_{\tilde{\mathcal{H}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id})$ (194)
$\leq\left(1+\frac{\lambda\sigma_{\mathrm{max}}(\nabla^{2}V)}{2}\right)\mathrm{d}_{\tilde{\mathcal{H}}_{\mu_{k}}}(\mathrm{T}_{k+1},\mathrm{Id}).$
@@ I.1 Three-point identity and inequality
$=\langle\nabla\phi(\mathrm{T}),\mathrm{S}-\mathrm{T}\rangle_{L^{2}(\mu)}-\langle\nabla\phi(\mathrm{U}),\mathrm{S}-\mathrm{T}\rangle_{L^{2}(\mu)}$
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathrm{T}_{0}\in L^{2}(\mu)$ and $\phi_{\mu}:L^{2}(\mu)\to\mathbb{R}$ convex, and Gâteaux differentiable. Let $\psi:L^{2}(\mu)\to\mathbb{R}$ be convex, proper and lower-semicontinuous. Assume there exists $\mathrm{T}^{*}=\operatorname*{argmin}_{\mathrm{T}\in L^{2}(\mu)}\ \mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{T}_{0})+\psi(\mathrm{T})$. Then, for all $\mathrm{T}\in L^{2}(\mu)$,
@@
Since $\phi$ and $\psi$ are proper, convex and lower-semicontinuous, and $\mathrm{T}\mapsto\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{T}_{0})$ is continuous (since $\phi_{\mu}$ is continuous), thus by [92, Theorem 3.30], $\partial\mathrm{J}(\mathrm{T}^{*})=\partial\psi(\mathrm{T}^{*})+\partial\mathrm{d}_{\phi_{\mu}}(\cdot,\mathrm{T}_{0})(\mathrm{T}^{*})$.
@@
$\psi(\mathrm{T})$ $\geq\psi(\mathrm{T}^{*})-\big{(}\langle\nabla\phi_{\mu}(\mathrm{T}^{*}),\mathrm{T}-\mathrm{T}^{*}\rangle_{L^{2}(\mu)}-\langle\nabla\phi_{\mu}(\mathrm{T}_{0}),\mathrm{T}-\mathrm{T}^{*}\rangle_{L^{2}(\mu)}\big{)}$ (198)
$=\psi(\mathrm{T}^{*})-\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{T}_{0})+\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{T}^{*})+\mathrm{d}_{\phi_{\mu}}(\mathrm{T}^{*},\mathrm{T}_{0}).$
@@
Actually we can restrict $\psi$ to be convex along $\big{(}(1-t)\mathrm{T}^{*}+t\mathrm{T}\big{)}_{\#}\mu$. In that case, $\mathrm{d}_{\psi}(\mathrm{T},\mathrm{T}^{*})=\psi(\mathrm{T})-\psi(\mathrm{T}^{*})-\langle\varphi,\mathrm{T}-\mathrm{T}^{*}\rangle_{L^{2}(\mu)}\geq 0$ for $\varphi\in\partial\psi(\mathrm{T}^{*})$ (by Proposition 11) and we still have $\partial\psi(\mathrm{T}^{*})+\partial\mathrm{d}_{\phi_{\mu}}(\cdot,\mathrm{T}_{0})(\mathrm{T}^{*})\subset\partial\mathrm{J}(\mathrm{T}^{*})$ (see [92, Theorem 3.30]) so that we can conclude.

### v2

@@ H.8 Proof of Lemma 11
On one hand, we have $\nabla_{2}\tilde{G}_{\mu}^{i}\big{(}\mathrm{T}_{s},\mathrm{T}_{s}(x)\big{)}=\nabla\frac{\partial}{\partial x_{i}}\frac{\delta\mathcal{F}}{\delta\mu}\big{(}(\mathrm{T}_{s})_{\#}\mu\big{)}\big{(}\mathrm{T}_{s}(x)\big{)}$. On the other hand, let us compute $\nabla_{1}\tilde{G}_{\mu}^{i}(\mathrm{T},x)$. First, we define the shorthands $\tilde{g}_{\mu}^{x,i}(\mathrm{T})=\tilde{G}_{\mu}^{i}(\mathrm{T},x)=\frac{\partial}{\partial x_{i}}\frac{\delta\mathcal{F}}{\delta\mu}(\mathrm{T}_{\#}\mu)(x)$ and $g^{x,i}(\nu)=\frac{\partial}{\partial x_{i}}\frac{\delta\mathcal{F}}{\delta\mu}(\nu)(x)$. Since $\tilde{g}_{\mu}^{x,i}(\mathrm{T})=g^{x,i}(\mathrm{T}_{\#}\mu)$, applying Proposition 1, we know that $\nabla_{1}\tilde{G}_{\mu}(\mathrm{T},x)=\nabla\tilde{g}_{\mu}^{x,i}(\mathrm{T})=\nabla_{\mathrm{W}_{2}}g^{x,i}(\mathrm{T}_{\#}\mu)\circ\mathrm{T}$.
@@ H.10 Proof of Proposition 13
    $\begin{cases}\langle\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{1}},\mathrm{T}_{t_{2}}-\mathrm{T}_{t_{1}}\rangle_{L^{2}(\mu)}\leq\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}-\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}\\ \langle\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{2}},\mathrm{T}_{t_{1}}-\mathrm{T}_{t_{2}}\rangle_{L^{2}(\mu)}\leq\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}-\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}.\end{cases}$ (168)
@@
    $\langle\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{2}}-\nabla_{W_{2}}\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}\circ\mathrm{T}_{t_{1}},\mathrm{T}_{t_{2}}-\mathrm{T}_{t_{1}}\rangle_{L^{2}(\mu)}\geq 0.$ (169)
@@
    (c4) $\implies$ (c1). Let $t_{1},t_{2}\in[0,1]$ and $\varphi(t)=\mathcal{F}(\tilde{\mu}_{t}^{t_{1}\to t_{2}})$ for all $t\in[0,1]$. From [125, Equation 16.5],
@@
    where $G$ is the Green function defined as $G(s,t)=s(1-t)\mathbbm{1}_{\{s\leq t\}}+t(1-s)\mathbbm{1}_{\{t\leq s\}}\geq 0$ [125, Equation 16.6]. Then, $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t})\geq 0$ implies that $\int_{0}^{1}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\varphi(s)G(s,t)\ \mathrm{d}s\geq 0$, and thus
    $\varphi(t)=\mathcal{F}(\tilde{\mu}_{t}^{t_{1}\to t_{2}})\leq(1-t)\varphi(0)+t\varphi(1)=(1-t)\mathcal{F}\big{(}(\mathrm{T}_{t_{1}})_{\#}\mu\big{)}+t\mathcal{F}\big{(}(\mathrm{T}_{t_{2}})_{\#}\mu\big{)}.\qed$ (174)
H.11 Proof of Proposition 24
@@
$\nabla\mathrm{J}(\tilde{\mathrm{T}}_{k+1})$ $=\nabla\phi_{\mu_{k}}(\tilde{\mathrm{T}}_{k+1})-\nabla\phi_{\mu_{k}}(\mathrm{Id})+\tau\big{(}\nabla_{\mathrm{W}_{2}}\mathcal{G}(\mu_{k})+\nabla_{\mathrm{W}_{2}}\mathcal{H}\big{(}(\tilde{\mathrm{T}}_{k+1})_{\#}\mu_{k}\big{)}\circ\tilde{\mathrm{T}}_{k+1}\big{)}$
@@
$=\nabla\phi_{\mu_{k}}(\tilde{\mathrm{T}}_{k+1})+\tau\nabla_{\mathrm{W}_{2}}\mathcal{H}\big{(}(\tilde{\mathrm{T}}_{k+1})_{\#}\mu_{k}\big{)}\circ\tilde{\mathrm{T}}_{k+1}-\nabla\phi_{\mu_{k}}(\mathrm{S}_{k+1}).$ (175)
@@
But $\mathrm{T}_{\phi_{\nu_{k+1}}}^{\nu_{k+1},\mu_{k+1}}\in L^{2}(\nu_{k+1})$, so this is a contradiction. So, we can conclude that the two schemes are equivalent, and moreover, $\tilde{\mathrm{T}}_{k+1}=\mathrm{T}_{\phi_{\nu_{k+1}}}^{\nu_{k+1},\mu_{k+1}}\circ\mathrm{S}_{k+1}$. ∎
H.12 Proof of Proposition 25
Let $\psi(\mathrm{T})=\tau\big{(}\langle\nabla_{\mathrm{W}_{2}}\mathcal{G}(\mu_{k}),\mathrm{T}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\mathcal{H}(\mathrm{T}_{\#}\mu_{k})\big{)}$. Since ~$\tilde{\mathcal{H}}_{\mu_{k}}$ is convex on $L^{2}(\mu_{k})$, $\psi$ is convex, and we can apply the three-point inequality (Lemma 29) and for all $\mathrm{T}\in L^{2}(\mu_{k})$,
@@ H.12 Proof of Proposition 25
$\mathcal{H}(\mu_{k+1})+\mathcal{G}(\mu_{k+1})$ $\leq\mathcal{H}(\mu_{k+1})+\mathcal{G}(\mu_{k})+\langle\nabla_{\mathrm{W}_{2}}\mathcal{G}(\mu_{k}),\tilde{\mathrm{T}}_{k+1}-\mathrm{Id}\rangle_{L^{2}(\mu_{k})}+\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\tilde{\mathrm{T}}_{k+1},\mathrm{Id})$
@@
$\quad-\frac{1}{\tau}\mathrm{d}_{\phi_{\mu_{k}}}(\mathrm{T},\tilde{\mathrm{T}}_{k+1}).$ (185)
@@
Plugging this into (H.12), we get
@@
Using $\mathrm{T}=\mathrm{Id}$ in (H.12), we observe that $\mathcal{F}(\mu_{k})\leq\mathcal{F}(\mu_{\ell})$ for all $\ell\leq k$. Moreover, $\mathrm{W}_{\phi}(\nu,\mu_{k})\geq 0$. Thus, applying Lemma 30 with $f=\mathcal{F}$, $c=\mathcal{F}(\nu)$ and $g=\mathrm{W}_{\phi}(\nu,\cdot)$, we obtain
$\forall k\geq 1,\ \mathcal{F}(\mu_{k})-\mathcal{F}(\nu)\leq\frac{\alpha}{\left(\frac{\frac{1}{\tau}}{\frac{1}{\tau}-\alpha}\right)^{k}-1}\mathrm{W}_{\phi}(\nu,\mu_{0})\leq\frac{\frac{1}{\tau}-\alpha}{k}\mathrm{W}_{\phi}(\nu,\mu_{0}).\qed$ (189)
H.13 Proof of Lemma 26
@@
and thus with $u$ being $(1-\tau\beta)$-convex relative to $V$ (for $\tau\beta\leq 1)$. For $\tau\beta<1$, it is equivalent to $u-(1-\tau\beta)V$ convex, i.e. $\langle\nabla u(x)-\nabla u(y),x-y\rangle\geq(1-\tau\beta)\langle\nabla V(x)-\nabla V(y),x-y\rangle\geq 0$. Since $V$ is strictly convex, $\nabla u$ is injective.
@@
To conclude, applying [5, Lemma 5.5.3], $\mathrm{T}_{\#}\mu$ is absolutely continuous with respect to the Lebesgue measure. ∎
H.14 Proof of Proposition 27
@@
Recall from (61) that $\mathrm{T}_{k+1}(x)=\big{(}(1-\tau)\Sigma_{k+1}\Sigma_{k}^{-1}+\tau\Sigma_{k+1}\Sigma^{-1}\big{)}x+cst$, thus $\nabla\mathrm{T}_{k+1}$ is a constant. Using the computations of [43, Appendix B.2],
@@
Moreover, by [33, Lemma 10], $\mu\mapsto\|\Sigma_{\mu}\|_{\mathrm{op}}$ is convex along generalized geodesics, and thus $\Sigma_{\mu_{t}}\preceq\lambda I_{d}$ and $\|\Sigma_{\mu_{t}}\|_{\mathrm{op}}\leq\lambda$ [43]. Hence, noting $\sigma_{\mathrm{max}}(M)$ the largest eigenvalue of some matrix $M$,
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{H}(\mu_{t})\geq\frac{1}{\lambda}\|\mathrm{T}_{k+1}-\mathrm{Id}\|_{L^{2}(\mu_{k})}^{2}$ $\geq\frac{1}{\lambda\sigma_{\mathrm{max}}(\nabla^{2}V)}\int(\mathrm{T}_{k+1}-\mathrm{Id})^{T}\nabla^{2}V(T_{k+1}-\mathrm{Id})\mathrm{d}\mu_{k}$
$=\frac{2}{\lambda\sigma_{\mathrm{max}}(\nabla^{2}V)}\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{V}(\mu_{t}).$ (198)
@@
[…]
@@ I.1 Three-point identity and inequality
$=\langle\nabla\phi(\mathrm{T}),\mathrm{S}-\mathrm{T}\rangle_{L^{2}(\mu)}-\langle\nabla\phi(\mathrm{U}),\mathrm{S}-\mathrm{T}\rangle_{L^{2}(\mu)}.$
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathrm{T}_{0}\in L^{2}(\mu)$ and $\phi_{\mu}:L^{2}(\mu)\to\mathbb{R}$ convex, and Gâteaux differentiable. Let $\psi:L^{2}(\mu)\to\mathbb{R}$ be convex, proper and lower semicontinuous. Assume there exists $\mathrm{T}^{*}=\operatorname*{argmin}_{\mathrm{T}\in L^{2}(\mu)}\ \mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{T}_{0})+\psi(\mathrm{T})$. Then, for all $\mathrm{T}\in L^{2}(\mu)$,
@@
Since $\phi$ and $\psi$ are proper, convex and lower semicontinuous, and $\mathrm{T}\mapsto\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{T}_{0})$ is continuous (since $\phi_{\mu}$ is continuous), thus by [99, Theorem 3.30], $\partial\mathrm{J}(\mathrm{T}^{*})=\partial\psi(\mathrm{T}^{*})+\partial\mathrm{d}_{\phi_{\mu}}(\cdot,\mathrm{T}_{0})(\mathrm{T}^{*})$.
@@
[…]
@@
Actually we can restrict $\psi$ to be convex along $\big{(}(1-t)\mathrm{T}^{*}+t\mathrm{T}\big{)}_{\#}\mu$. In that case, $\mathrm{d}_{\psi}(\mathrm{T},\mathrm{T}^{*})=\psi(\mathrm{T})-\psi(\mathrm{T}^{*})-\langle\varphi,\mathrm{T}-\mathrm{T}^{*}\rangle_{L^{2}(\mu)}\geq 0$ for $\varphi\in\partial\psi(\mathrm{T}^{*})$ (by Proposition 13) and we still have $\partial\psi(\mathrm{T}^{*})+\partial\mathrm{d}_{\phi_{\mu}}(\cdot,\mathrm{T}_{0})(\mathrm{T}^{*})\subset\partial\mathrm{J}(\mathrm{T}^{*})$ (see [99, Theorem 3.30]) so that we can conclude.